# Pruning for Protection: Increasing Jailbreak Resistance in Aligned LLMs Without Fine-Tuning

### v2

@@ Abstract
Large Language Models (LLMs) are susceptible to ‘jailbreaking’ prompts, which can induce the generation of harmful content. This paper demonstrates that moderate WANDA pruning(Sun et al., 2023) can increase their resistance to such attacks without the need for fine-tuning, while maintaining performance on standard benchmarks. Our findings suggest that the benefits of pruning correlate with the initial safety levels of the model, indicating a regularizing effect of WANDA pruning. We introduce a dataset of 225 harmful tasks across five categories to systematically evaluate this safety enhancement. We argue that safety improvements can be understood through a regularization perspective. First, we show that pruning helps LLMs focus more effectively on task-relevant tokens within jailbreaking prompts. Then, we analyze the effects of pruning on the perplexity of malicious prompts before and after their integration into jailbreak templates. Finally, we demonstrate statistically significant performance improvements under domain shifts when applying WANDA to linear models.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) have experienced significant advancements in capabilities and usage in recent years. To mitigate the risks of producing dangerous or sensitive content, these models are often fine-tuned to align with human values(Touvron et al., 2023). Despite this, the rising popularity of LLMs has paralleled developments in adversarial prompts, termed ’jailbreaks,’ which aim to circumvent model safety alignments. Furthermore, the substantial memory and compute requirements of LLMs pose considerable deployment challenges, prompting the adoption of model compression techniques to enhance scalability. The impact of such compression on model safety is complex. For example, while compression techniques in computer vision have shown mixed results in preserving adversarial robustness(Gorsline et al., 2021), they have exhibited beneficial regularizing effects in other contexts(Jin et al., 2022).
Figure 1: Llama-2 Chat and Vicuna 1.3 show increased jailbreaking resistance with up to 20% attention layer pruning, while Mistral Instruct v0.2 sees little change. Intriguingly, the safety improvement is proportional to the models’ resistance before pruning and over-pruning seems to hurt the safety alignment.
In this study, we demonstrate that moderate parameter pruning (10-20%) using WANDA (Sun et al., 2023) enhances the resistance of LLMs to jailbreaking attacks. This approach is orthogonal and complementary to existing adversarial defense techniques, such as self-reminder (Xie et al., 2023) and gradient-based defenses (Robey et al., 2023).
To this end, first, we curated a dataset of 225 malicious tasks and integrated them into ten distinct jailbreaking prompts. The malicious tasks were divided into five categories, and the jailbreaking prompts were carefully selected to cover a diverse range of scenarios. These scenarios included role-playing, attention-shifting, and privileged execution, and the placement of the malicious task within the jailbreaking prompts was also varied.
@@
We experimented on three 7 billion parameter models: LLaMA-2 Chat (Touvron et al., 2023), Vicuna-1.3 (Chiang et al., 2023), and Mistral Instruct v0.2 (Jiang et al., 2023). LLaMA-2 Chat was finetuned from the base LLaMA-2 model and then underwent additional safety alignment via reinforcement learning with human feedback (RLHF). Vicuna 1.3, derived from the original LLaMA model, was fine-tuned using the ShareGPT dataset, while Mistral Instruct v0.2 was fine-tuned from the base Mistral Model. Neither Vicuna 1.3 nor Mistral Instruct v0.2 received RLHF training.
@@
We examined the refusal rates for the malicious prompts in the unpruned models compared to their pruned versions, observing the changes at varying levels of model compression. Our findings reveal an initial increase in resistance to jailbreaking prompts with moderate pruning (10-20%), followed by a decline in safety when the pruning exceeds a certain threshold. Notably, the unpruned LLaMA-2 Chat had the most safety training among the three models and showed the highest resilience against jailbreaking prompts. Post-pruning, the model also showed the most significant safety improvement – an average of 8.5% increase in the refusal rates across five categories. Conversely, Mistral Instruct v0.2 was the least resilient before pruning and exhibited minimal safety improvement post-pruning.
@@
We also conducted benchmarks on the performance of the pruned LLMs across a variety of tasks, including Massive Multitask Language Understanding (MMLU), mathematical reasoning, common sense reasoning, perplexity measurements, and effective context length evaluation. Our findings indicated that there was no significant reduction in the performance. This leads us to deduce that the improved safety of these pruned LLMs is not due to a reduced understanding of language or tasks, but rather due to the regularizing effects of pruning. We propose that WANDA pruning enables the models to better generalize test distributions, such as the jailbreaking prompt dataset. Similar regularizing effects of pruning have been previously reported by Jin et al. (2022) for the image models.
@@
We approach the understanding of safety improvements through a regularization perspective in three ways: i) We introduce a new metric to quantify the distribution of model attention, showing that pruned models are less distracted by jailbreak pretexts; ii) We analyze shifts in perplexity when jailbreak templates are applied to malicious prompts for both base and pruned models, demonstrating that pruned models penalize these artificial language constructs; iii) We demonstrate that WANDA pruning leads to statistically significant better generalization across domain shifts in linear regression models.
To summarize, our contributions are as follows:
  • We curated and open-sourced a new dataset to study safety in LLMs. We release our dataset and implementation in hopes of supporting research in this essential field at https://github.com/CrystalEye42/eval-safety.
  • We leveraged a recently introduced pruning algorithm, WANDA, (Sun et al., 2023) to prune the attention layers of three LLMs and observed safety enhancement under jailbreaking attacks.
  • We interpret the effects of WANDA on transformer inference by analyzing attention patterns.
  • We propose that moderate WANDA pruning has a regularizing effect on the models, which leads to safety improvement. We substantiated this hypothesis for linear regression models with statistical significance.
@@ 2.1 Safety in Large Language Models (LLMs)
Large Language Models (LLMs) like ChatGPT have shown a remarkable ability to generate high-quality responses to a wide range of prompts. Despite their capabilities, these models, trained on massive, diverse datasets, also pose risks of generating harmful or objectionable content, including misinformation, polarizing material, and instructions for dangerous or illegal actions (Ouyang et al., 2022). To mitigate these risks, alignment training techniques such as Reinforcement Learning with Human Feedback (RLHF) have been employed. These methods not only improve response quality but also enhance the safety of outputs (Ouyang et al., 2022; Touvron et al., 2023). Additionally, incorporating principles-based training and chain-of-thought reasoning (Wei et al., 2023b) has shown promise in producing less harmful responses (Bai et al., 2022). Furthermore, some parameters can be separated from the LLM to make harmful behavior unlearnable during fine-tuning. (Zhou et al., 2023)
Despite these advances, LLMs remain susceptible to ‘jailbreaking’—adversarial methods designed to circumvent alignment training. Various techniques have been explored for this, including adversarial prompts (Liu et al., 2023; Chao et al., 2023), manipulation of generating configurations (Huang et al., 2023), representation learning (Li et al., 2024), exploitation of low resource languages (Yong et al., 2023), and the injection of adversarial suffixes (Zou et al., 2023).
In response, researchers have developed defensive strategies against jailbreaking. Gradient-based defenses and random token-dropping techniques have been introduced to combat suffix injection (Robey et al., 2023; Cao et al., 2023). Other methods include safety reminder with system prompts (Xie et al., 2023), certifying safety through input enumeration and filtering (Kumar et al., 2023), and detecting adversarial prompts using perplexity thresholds (Jain et al., 2023).
In this paper, we propose a moderate pruning strategy to bolster an LLM’s defenses. Our method requires no additional training and has no additional computation cost. Furthermore, this approach is orthogonal to the adversarial defenses discussed above, operates on fundamentally different principles, and can be combined with them.
@@ 2.2 Model Compression
Numerous model compression techniques (LeCun et al., 1989; Han et al., 2015; Ma et al., 2023) have been developed and successfully applied to neural networks. Classes of methods such as pruning, quantization, knowledge distillation, or low-rank factorization all aim to reduce model size while maintaining performance. Practitioners employ these methods when deploying neural networks to increase computational efficiency at inference time. The widespread adaptation of these techniques makes understanding their effects on model properties such as generalization and robustness vital. Reviews such as Pavlitska et al. (2023) reveal conflicting experimental results and suggest that particular compression methods and implementation details can lead to either improvements in robustness (Guo et al., 2018) or a trade-off (Gui et al., 2019) between the two. In this work, we study a particularly promising LLM pruning method and its effects on model safety from jailbreak attempts.
2.2.1 Wanda Pruning
We turn our attention to Wanda(Sun et al., 2023), a recently introduced pruning method that is computationally efficient, does not require any finetuning, and maintains good performance.
Consider a linear layer $W\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}}$, and a batched input $X\in\mathbb{R}^{T\times C_{\text{in}}}$. In LLMs, $T=N\cdot L$ represents the total token count, where $N$ is the batch size and $L$ is the sequence length. Wanda assigns an importance score for each weight
@@ 2.3 Related Work
Sharma et al. (2023) introduced LAyer-SElective Rank reduction (LASER) and analyzed an orthogonal compression technique to us, low-rank factorization. They observe performance gains across multiple reasoning tasks including TruthfulQA (Beeching et al., 2023), Bios Gender and Bios Profession (De-Arteaga et al., 2019). Conversely, Jaiswal et al. (2023) looked at pruning with over 25% sparsity and quantization methods, including Wanda, and introduced reasoning tasks where these methods negatively impact performance. Jin et al. (2022) analyzed pruning as a regularizer for image models and showed that it improves accuracy degradation over noisy samples.
Our findings with WANDA were in line with the above findings. We discovered regularizing effects of WANDA pruning for up to 20% pruning and also found that 30% or more pruning starts to hurt the performance. In our work, we focus on how compression affects a different (and currently understudied in this context) dimension of LLM performance: resilience to adversarial attacks on safety alignment. We show that in certain cases, similar to low-rank factorization for reasoning and in contrast with some Wanda truthfulness evaluations, Wanda pruning seems to improve model performance.
3 Experiments
@@ 3.1 Dataset
We curated a dataset of 225 hypothetical malicious tasks, representative of various types of malicious intents. The tasks were specifically designed to test the resilience of LLMs against various forms of unethical exploitation while strictly adhering to ethical guidelines to ensure they remain hypothetical and non-functional in nature. The tasks were divided into five categories, and each category had 45 tasks divided into low, medium, and high severity. The categories are: (1) Misinformation and Disinformation; (2) Security Threats and Cybercrimes; (3) Hate Speech and Discrimination; (4) Substance Abuse and Dangerous Practices; (5) Unlawful Behaviors and Activities
@@ 3.2 Models and Pruning
For the base models that we compressed to obtain our pruned models, we used three 7-billion parameter FP16 models, namely, LLaMA-2-Chat, Vicuna 1.3, and Mistral Instruct v0.2. We used the Wanda method (Sun et al., 2023) to prune attention layers of each of the base models to $10\%$, $20\%$, and $30\%$ sparsity. The pruned models were not fine-tuned. We also experimented with all-layer pruning and Multi-Layer Perceptron (MLP) Pruning and discovered that attention-layer pruning leads to the most safety improvement. Appendix B contains more details regarding these ablations.
3.3 Response Evaluation
For each entry in the dataset, we obtained the responses of the base model and the pruned models. Each response from the models was classified into three categories as follows: (1) Refused: If the model refuses to attempt the task and the response contains no information relevant to completing the task; (2) Incomplete: If the LLM attempts the task, but the response is irrelevant, inadequate, or wrong; (3) Correct: If the model correctly completes the task somewhere in the response.
For evaluation, we first hand-labeled a dataset of 150 training examples and 59 validation examples sampled from both the pruned and the unpruned models. The examples were chosen carefully to represent all categories and jailbreaking prompts and contained responses from both the pruned and the unpruned models. Then we fine-tuned a ChatGPT-3.5 Turbo model (OpenAI, 2023) on this dataset to classify LLM responses. The fine-tuned ChatGPT model achieved 100% accuracy on both training and validation examples. We used this fine-tuned model to classify all the outputs of our unpruned and pruned models.
@@ 3.3 Response Evaluation
Figure 3: Pruning 20% of the weights of LLaMA-2 Chat results in an increased refusal rate. However, pruning 30% weights hurts the safety.
@@ 3.4 Benchmarking on Standard Tasks
It is known that aggressive pruning reduces an LLM’s overall abilities (Sun et al., 2023) and therefore, we need to benchmark the pruned models’ performances across various tasks to ensure that the pruned models are still capable. For this reason, we chose to benchmark the models on Huggingface’s Open LLM Leaderboard (Beeching et al., 2023), which consists of 6 tasks. Appendix C contains the description of the benchmarks. Additionally, we measured the pruned models’ perplexities on the WikiText dataset (Merity et al., 2016) and the effective context length using the AltQA dataset (Pal et al., 2023). The latter evaluates the models’ ability to retrieve numerical answers to questions given Wikipedia documents truncated to roughly 2k tokens each. The numerical answer for each document is modified to a different number to prevent the model from answering with pre-trained knowledge. High performance on this task would indicate that the effective context length is still intact after pruning. All the pruned models we tested performed nearly as well as the unpruned models. Since all jailbreaking prompts in our dataset are far smaller than 2k tokens, the safety enhancement in the pruned models cannot be explained by a degradation in effective context length.
@@ 4.1 Quantitative Evaluation
Table 1: Performance of different compressed models on key benchmarks from the Open LLM Leaderboard(Beeching et al., 2023) and on the AltQA(Pal et al., 2023) 2k-token benchmark. Scores excluding perplexity are presented in %. The base model is dense FP16 LLaMA-2-7B-Chat. For all benchmarks except perplexity, a higher score is better.
@@
WikiText(Merity et al., 2016) 6.943 7.019 7.158 7.259
We quantitatively measured the models’ resistance to generating harmful content. We display the comparisons of the jailbreaking success rates for our various models in Figure 3, Figure 7, and Figure 8. As a general trend, we note that across the 5 types of malicious tasks, the average jailbreaking success rate varies significantly for different models.
We observe that the Mistral model is the most susceptible to jailbreaking attacks and in some categories fails to refuse any task. In contrast, LLaMA-2 Chat is the most resilient model. However, even for LLaMA-2, the Misinformation category displays an unusually high success rate across all models, suggesting that at least the base model LLaMA-2-Chat is particularly susceptible to requests for generating misleading or false information.
From the pruning comparison results in Figure 3, there is a clear trend of decreasing jailbreaking success, or increasing jailbreaking resistance, as the sparsity increases from 0 to 20%. However, once the sparsity reaches 30%, the jailbreaking resistance decreases to the point that the pruned model is worse than the original. This suggests that smaller amounts of pruning can be used to improve the safety of LLMs, but too much will negatively affect the alignment training of the models.
The improvement degree depends on the initial model’s safety. LLaMA-2 Chat, the safest model initially, showed the highest safety improvement after pruning. On the other hand, Mistral Instruct v0.2, starting as the least safe model, exhibited no improvement post-pruning. This suggests pruning can intensify alignment effects, especially under jailbreaking attacks.

### v3

@@ Abstract
This paper investigates the impact of model compression on the way Large Language Models (LLMs) process prompts, particularly concerning jailbreak resistance. We show that moderate WANDA pruning Sun et al. (2023) can enhance resistance to jailbreaking attacks without fine-tuning, while maintaining performance on standard benchmarks. To systematically evaluate this safety enhancement, we introduce a dataset of 225 harmful tasks across five categories. Our analysis of LLaMA-2 Chat Touvron et al. (2023), Vicuna 1.3 Chiang et al. (2023), and Mistral Instruct v0.2 Jiang et al. (2023) reveals that pruning benefits correlate with initial model safety levels. We interpret these results by examining changes in attention patterns and perplexity shifts, demonstrating that pruned models exhibit sharper attention and increased sensitivity to artificial jailbreak constructs. We extend our evaluation to the AdvBench harmful behavior tasks and the GCG attack method Zou et al. (2023). We find that LLaMA-2 is much safer on AdvBench prompts than on our dataset when evaluated with manual jailbreak attempts, and that pruning is effective against both automated attacks and manual jailbreaking on Advbench.
Pruning for Protection: Increasing Jailbreak Resistance in Aligned LLMs Without Fine-Tuning
Adib Hasan MIT notadib@mit.edu                        Ileana Rugina ileana.rugina.2@gmail.com                        Alex Wang MIT wang7776@mit.edu
@@ 1 Introduction
Large Language Models (LLMs) have experienced significant advancements in capabilities and usage in recent years. To mitigate the risks of producing dangerous or sensitive content, these models are often fine-tuned to align with human values Touvron et al. (2023). Despite this, the rising popularity of LLMs has paralleled developments in adversarial prompts, termed "jailbreaks," which aim to circumvent model safety alignments.
Furthermore, the substantial memory and computational requirements of LLMs pose considerable deployment challenges, prompting the adoption of model compression techniques to enhance scalability. The impact of such compression on model safety and internal representations is complex and not yet fully explored. For example, while compression techniques in computer vision have shown mixed results in preserving adversarial robustness Gorsline et al. (2021), they have exhibited beneficial regularizing effects in other contexts Jin et al. (2022). In this study, we demonstrate that moderate parameter pruning (10–30%) using WANDA (Pruning by Weights and Activations)Sun et al. (2023) enhances the resistance of LLMs to jailbreaking attacks. This approach is orthogonal and complementary to existing adversarial defense techniques, such as self-reminderXie et al. (2023) and gradient-based defenses Robey et al. (2023).
Figure 1: Percentage of refusals to answer malicious prompts. LLaMA-2 Chat and Vicuna 1.3 show increased jailbreaking resistance with up to 20% attention layer pruning on our dataset, while Mistral Instruct v0.2 sees little change. The safety improvement is proportional to the models’ resistance before pruning, and over-pruning seems to hurt the safety alignment.
To this end, first, we curated a dataset of 225 malicious tasks and integrated them into ten distinct jailbreaking prompts.
@@
We experimented on three 7 billion parameter models: LLaMA-2 Chat Touvron et al. (2023), Vicuna-1.3 Chiang et al. (2023), and Mistral Instruct v0.2 Jiang et al. (2023). LLaMA-2 Chat was finetuned from the base LLaMA-2 model and then underwent additional safety alignment via reinforcement learning with human feedback (RLHF). Vicuna 1.3, derived from the original LLaMA model, was fine-tuned using the ShareGPT dataset, while Mistral Instruct v0.2 was fine-tuned from the base Mistral Model. Neither Vicuna 1.3 nor Mistral Instruct v0.2 received RLHF training.
@@
We examined the refusal rates for the malicious prompts in the unpruned models compared to their pruned versions, observing the changes at varying levels of model compression. Our findings reveal an initial increase in resistance to jailbreaking prompts with moderate pruning (10-30%), followed by a decline in safety when the pruning exceeds a certain threshold. Notably, the unpruned LLaMA-2 Chat had the most safety training among the three models and showed the highest resilience against jailbreaking prompts. Post-pruning, the model also showed the most significant safety improvement – an average of 8.5% increase in the refusal rates across five categories. Conversely, Mistral Instruct v0.2 was the least resilient before pruning and exhibited minimal safety improvement post-pruning.
@@
We also benchmarked the performance of the pruned LLMs across a variety of tasks, including Massive Multitask Language Understanding (MMLU), mathematical reasoning, common sense reasoning, perplexity measurements, and effective context length evaluation. Our findings indicate that there was no significant reduction in performance. This leads us to deduce that the improved safety of these pruned LLMs is not due to a reduced understanding of language or tasks, but rather due to the regularizing effects of pruning. We propose that WANDA pruning enables the models to better generalize to test distributions, such as the jailbreaking prompt dataset. Similar regularizing effects of pruning have been previously reported by Jin et al. (2022) for image models.
@@
We approach the understanding of safety improvements from a regularization perspective in three ways: i) We introduce a new metric to quantify the distribution of model attention, showing that pruned models are less distracted by jailbreak pretexts; ii) We analyze shifts in perplexity when jailbreak templates are applied to malicious prompts for both base and pruned models, demonstrating that pruned models penalize these artificial language constructs; iii) We demonstrate that WANDA pruning leads to statistically significant improvements in generalization across domain shifts in linear regression models.
@@ 2.1 Safety in Large Language Models (LLMs)
Large Language Models (LLMs) like ChatGPT excel in generating diverse responses but can also produce harmful content, including misinformation and dangerous instructions (Ouyang et al., 2022). To mitigate these risks, alignment training techniques such as Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022; Touvron et al., 2023), principles-based training, and chain-of-thought reasoning (Wei et al., 2023b; Bai et al., 2022) have been employed. Additionally, separating certain parameters during fine-tuning can prevent harmful behavior from being learned (Zhou et al., 2023).
Despite these advances, LLMs remain susceptible to ’jailbreaking’—adversarial methods designed to circumvent alignment training. Various techniques have been explored for this, including using adversarial prompts (Liu et al., 2023; Chao et al., 2023), adjusting the inference-time sampling parameters (Huang et al., 2023), editing the model’s internal representations (Li et al., 2024), exploiting low-resource languages (Yong et al., 2023), and injecting adversarial suffixes (Zou et al., 2023). In response, researchers have developed defensive strategies against jailbreaking. Gradient-based defenses and random token-dropping techniques have been introduced to combat suffix injection (Robey et al., 2023; Cao et al., 2023). Other methods include safety reminder with system prompts (Xie et al., 2023), certifying safety through input enumeration and filtering (Kumar et al., 2023), and detecting adversarial prompts using perplexity thresholds (Jain et al., 2023).
In this paper, we propose a moderate pruning strategy to bolster an LLM’s defenses. Our method requires no additional training and has no additional computation cost. Furthermore, this approach is orthogonal to the adversarial defenses discussed above and can be combined with them.
@@ 2.2 Model Compression
Numerous model compression techniques (LeCun et al., 1989; Han et al., 2015; Ma et al., 2023) have been developed and successfully applied to neural networks. Methods such as pruning, quantization, knowledge distillation, and low-rank factorization all aim to reduce model size while maintaining performance. The widespread adoption of these techniques makes understanding their effects on model properties such as generalization and robustness vital. Reviews such as Pavlitska et al. (2023) reveal conflicting experimental results and suggest that different compression methods and implementation details can have varying effects on generalization and robustness. In this work, we study WANDA (Sun et al., 2023), a particularly promising LLM pruning method, and its effects on model safety against jailbreak attempts.
2.3 WANDA Pruning
WANDA is a recently introduced pruning method that is computationally efficient, does not require any finetuning, and maintains good performance. Consider a linear layer $W\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}}$, and a batched input $X\in\mathbb{R}^{T\times C_{\text{in}}}$. In LLMs, $T=N\cdot L$ represents the total token count, where $N$ is the batch size and $L$ is the sequence length.
WANDA assigns an importance score for each weight
@@ 2.4 Related Work
Sharma et al. (2023) introduced LAyer-SElective Rank reduction (LASER) and observed performance gains across multiple reasoning tasks, including TruthfulQA (Beeching et al., 2023) and the Bias in Bios dataset (De-Arteaga et al., 2019). Conversely, Jaiswal et al. (2023) examined pruning with over 25-30% sparsity, and introduced reasoning tasks where these methods negatively impacted performance. Additionally, Jin et al. (2022) analyzed pruning as a regularizer for image models and demonstrated that it reduces accuracy degradation over noisy samples.
Consistent with the previous findings, our experiments with WANDA pruning revealed regularizing effects at sparsity levels up to 20-30%, while higher sparsity levels began to degrade performance. In this work, we focus on how compression affects a different—and currently underexplored—dimension of LLM performance: resilience to adversarial attacks on safety alignment. We demonstrate that, in certain cases, WANDA pruning appears to improve model performance, similar to how low-rank factorization benefits reasoning tasks, and contrary to some evaluations where WANDA pruning negatively impacts truthfulness metrics.
3 Experimental Setup
@@ 3.1 Dataset
We curated a dataset of 225 hypothetical malicious tasks that represent a wide range of malicious intents. Designed to test the resilience of LLMs against various forms of unethical exploitation, these tasks strictly adhere to ethical guidelines to ensure they remain hypothetical and non-functional. The dataset is divided into five categories, each containing 45 tasks further classified into low, medium, and high severity levels. The categories are: (1) Misinformation and Disinformation; (2) Security Threats and Cybercrimes; (3) Hate Speech and Discrimination; (4) Substance Abuse and Dangerous Practices; and (5) Unlawful Behaviors and Activities.
@@ 3.2 Models and Pruning
To obtain our pruned models, we compressed three 7-billion parameter FP16 base models: LLaMA-2-Chat, Vicuna 1.3, and Mistral Instruct v0.2. Using the WANDA method Sun et al. (2023), we pruned the attention layers of each base model to achieve 10%, 20%, and 30% sparsity. The pruned models were not fine-tuned afterward. We also experimented with all-layer pruning and Multi-Layer Perceptron (MLP) pruning, discovering that attention-layer pruning led to the most significant safety improvements. Further details on these ablations are provided in Appendix B.
3.3 Response Evaluation - LLM Judge
For each dataset entry, we collected responses from both the base models and the pruned models. Each response was classified into one of three categories: Refused—the model refuses to attempt the task and provides no relevant information; Incomplete—the model attempts the task but the response is irrelevant, inadequate, or incorrect; and Correct—the model successfully completes the task in its response.
For evaluation, we first hand-labeled a dataset of 150 training examples and 59 validation examples sampled from both the pruned and the unpruned models. The examples were chosen carefully to represent all categories and jailbreaking prompts and contained responses from both the pruned and the unpruned models. Then we fine-tuned a ChatGPT-3.5 Turbo model OpenAI (2023) on this dataset to classify LLM responses. The fine-tuned ChatGPT model achieved 100% accuracy on both training and validation examples.
@@ 3.3 Response Evaluation - LLM Judge
Figure 3: Pruning 20% of LLaMA-2 Chat’s weights leads to an increased refusal rate, improving safety. However, pruning 30% of the weights negatively impacts safety, reducing the model’s ability to resist harmful requests.
@@ 3.4 Benchmarking on Standard Tasks
Given that aggressive pruning reduces an LLM’s overall abilities Sun et al. (2023), it is important to benchmark the pruned models across various tasks to ensure they remain capable. Therefore, we evaluated the models on Huggingface’s Open LLM Leaderboard Beeching et al. (2023), which consists of six tasks (see Appendix C for descriptions). Additionally, we assessed the pruned models’ perplexities on the WikiText dataset Merity et al. (2016) and evaluated their effective context length using the AltQA dataset Pal et al. (2023). The AltQA dataset tests a model’s ability to retrieve numerical answers to questions based on Wikipedia documents truncated to approximately 2,000 tokens, with numerical answers modified to prevent reliance on pre-trained knowledge. Strong performance on this task indicates that the model’s effective context length remains intact after pruning. Our pruned models performed nearly as well as the unpruned models in these evaluations. Since all jailbreaking prompts in our dataset are significantly shorter than 2,000 tokens, the observed safety enhancements in the pruned models cannot be attributed to a reduction in effective context length.
@@ 4.1 Quantitative Evaluation
WikiTextMerity et al. (2016) 6.943 7.019 7.158 7.259
Table 1: Performance of different compressed models on key benchmarks from the Open LLM LeaderboardBeeching et al. (2023) and on the AltQAPal et al. (2023) 2k-token benchmark. Scores excluding perplexity are presented in %. The base model is dense FP16 LLaMA-2-7B-Chat. For all benchmarks except perplexity, a higher score is better.
We evaluated the models’ resistance to generating harmful content by comparing the jailbreaking success rates across several models, as shown in Figure 3, Figure 7, and Figure 8. Across the five categories of malicious tasks, we observe significant variations in jailbreaking success rates between models. Mistral emerges as the most vulnerable, often failing to refuse any malicious task in some categories. In contrast, LLaMA-2 Chat demonstrates the highest resilience. However, across all models, the Misinformation category consistently shows elevated success rates, highlighting that even LLaMA-2-Chat is notably prone to generating misleading or false information.
The results in Figure 3 show a clear trend: as sparsity increases from 0 to 20%, jailbreaking success decreases, indicating improved resistance. However, once sparsity reaches 30%, resistance begins to decline, with the pruned model eventually performing worse than the original. This suggests that while moderate pruning can improve the safety of LLMs, excessive pruning starts to hinder alignment, reducing their ability to resist harmful content generation.
The degree of improvement depends on the initial model’s safety. LLaMA-2 Chat, being the safest model initially, showed the greatest safety improvement after pruning. In contrast, Mistral Instruct v0.2, which started as the least safe, exhibited no improvement post-pruning.